Find the area under the standard normal curve than z=0.5


1
Expert's answer
2021-05-10T17:16:36-0400

So "P(Z>0.5)=P(0<z<\infty)-P(0<z<0.5)=0.5-0.19146=0.30854"


Area under normal curve of (z>0.5) is 0.30854.